examine the data in appendix of chap 19 locate the nominal and real gdp. calculate gdp deflator what are the growth of rate of real gdp and nominal gdp for 2007? what are the rate of inflation for 2007?Compare the rate of inflation using the GDP deflator with that of using CPI.

Book:Paul Samuelson 19th edition.
1
Expert's answer
2021-09-07T09:44:11-0400

"GDP \space deflator=\frac{nominal GDP}{real \space GDP}\times 100"


"2007\\=\frac{12000}{10000}\times100=120\%\\2006\\=\frac{10000}{8000}\times100=125\%"


rate of growth of

i)nominal GDP in 2007

"=\frac{nominal\space GDP\space in\space 2007-nominal\space GDP\space in\space 2006}{nominal \space GDP\space in\space 2006}\times 100\\=\frac{12000-10000}{10000}\times 100\\=20\%"


ii)real GDP in 2007

"=\frac{real\space GDP\space in\space 2007-real\space GDP\space in\space 2006}{real \space GDP\space in\space 2006}\times 100\\=\frac{10000-8000}{8000}\times 100\\=25\%"


rate of inflation in 2007

"\frac{deflator\space in\space 2006-deflator \space in \space 2006}{deflator\space in\space 2006}\times 100\\=\frac{120-125}{125}\times 100\\=-4\%"

in GDP deflator, inflation measures level of prices of purchase by government, consumers and business while in CPI the inflation measures level of prices of purchase by consumers only.